OpenStudy (anonymous):

How many solutions does this system of equations have? x-4y=13 5x-20y=63

OpenStudy (callisto):

for the 2nd equation, take out 5 as the common factor on the left hand side , and divide both sides by 5, can you show me what would you get?

OpenStudy (anonymous):

would it have none?

OpenStudy (callisto):

if you take out the common factor 5x-20y=63 5(x-4y) = 63 then divide both sides by 5 5(x-4y) /5= 63/5 (x-4) = 63/5

OpenStudy (anonymous):

so it would be infinite??

OpenStudy (anonymous):

it would only have 1.Correct??

OpenStudy (callisto):

after that , rewrite the 2 equations x-4y=13 => x-4y-13 =0 (x-4y) = 63/5 => x-4y -63/5 =0 So , you can see that they have the same slope, but different y-int So they are //lines, like this  So how many intercept(s) can you see?

OpenStudy (anonymous):

1

OpenStudy (callisto):

where?

OpenStudy (anonymous):

What i the answer??

OpenStudy (anonymous):

is*

OpenStudy (callisto):

sorry i won't give you the answer directly this time.. you should work it out yourself

OpenStudy (callisto):

have you ever seen 2 parallel lines having intercept(s) ?